Solve for x
x-a=-x+a+4b

Answer by unlockmath(1688)

Hello,
For this one x-a=-x+a+4b we first want to add 1x to both sides to get:
2x-a=a+4b
Now add 1a to both sides to result in:
2x=2a+4b
Divide both sides by 2 and we get:
x=a+2b
Make sense?
RJ
www.math-unlock.com